# Approximating Maximum Edge 2-Coloring

### v2

@@ 3.2 The proof of Lemma 1
Proof
Since $K$ is a plain cactus and we cannot apply Modification $3$, there have to be edges not allowed in a simple cactus. These cannot be degree-$4$ vertices, due to Lemma 14. Also, by Lemma 15, edges between vertices from $K$ cannot violate the properties of simple cacti. Thus there has to be a degree-$2$ vertex $w$ of $K$ which has a set $A$ of at least two adjacent vertices in $V(G)\setminus V(K)$. By Invariant 8, there is a vertex $w^{\prime}\in A$ with $ww^{\prime}\in E(H)$. Let $c$ be the color of the triangle containing $w$ and $c^{\prime}:=\chi(ww^{\prime})$.
We claim that there is no vertex $\hat{w}\in A$ with $\chi(w\hat{w})\neq c^{\prime}$. Since $w$ only sees two colors, the only possibility would be $\chi(w\hat{w})=c$. However, since $w$ sees two colors, $\hat{w}$ cannot be an inner vertex and by Lemma 13 it also cannot be a free vertex or an end vertex.
Next, we show that $\iota(w^{\prime})=w^{\prime\prime}$. Suppose $\iota(w^{\prime})\neq w^{\prime\prime}$. Then, by Lemma 13, it sees two colors which are distinct from $c$, and also distinct from $c^{\prime}$ since $ww^{\prime}\in E(H)$ and $w,w^{\prime}\notin\iota^{-1}(w^{\prime\prime})$. However, $w^{\prime\prime}$ also sees $c^{\prime}$ and therefore more than two colors, a contradiction. Thus $\iota(w^{\prime})=w^{\prime\prime}$. In particular, $\chi(ww^{\prime})=\chi(ww^{\prime\prime})=\chi(w^{\prime}w^{\prime\prime})$. ∎
With Lemma 16 we know that adding the triangle $\{w,w^{\prime},w^{\prime\prime}\}$ does not violate the conditions of plain cacti, i.e., adding it still satisfies Invariant 1. Since $\iota(w^{\prime})=w^{\prime\prime}$, also Invariants 2 and 3 are satisfied. Invariant 4 follows directly from Lemma 16. Since $ww^{\prime}\in E(H)$ by definition (within Algorithm 2), Invariant 5 follows. Invariant 6 follows by noting that $w\in\text{scope}(v)$, the edge $ww^{\prime}$ satisfies the conditions of the lemma, and thus the color class $c$ is added to the scope. Invariant 7 is a direct consequence of Invariant 6. Invariant 8 requires additional arguments.
Lemma 17
@@
Proof
Let $\tilde{u}$ be a degree-$2$ vertex. We distinguish the type of $\tilde{u}$. Due to Invariant 7, $\tilde{u}$ is not free. If $\tilde{u}$ is an inner vertex, only one of the two incident edges from $E(H)$ can be in $E(K)$, due to Invariant 4. Thus due to the other edge, the invariant is satisfied. If $\tilde{u}$ is an end vertex, it has to have an incident edge $\tilde{u}u^{\prime}$ not in $E(K)$, due to Modification $2$. If $u^{\prime}\in V(K)$, Lemma 15 implies the invariant. We therefore may assume that $u^{\prime}\notin V(K)$.
By Invariant 3, there is a vertex $\bar{u}\in V(K)$ with $\iota(\bar{u})=\tilde{u}$. In particular, $\bar{u}$ is an inner vertex, $\bar{u}\tilde{u}\in E(K)$ and $\bar{u}\tilde{u}\notin E(H)$. Let $\hat{u}$ be the third vertex in the same triangle. By Invariant 4, $\chi(\hat{u}\tilde{u})=\chi(\hat{u}\bar{u})=\chi(\tilde{u}\bar{u})$. But then $\hat{u}\bar{u}\in E(H)$ as otherwise $\bar{u}$ would see three colors. Thus $\hat{u}\tilde{u}\notin E(H)$. However, since $u$ is an end vertex, it is incident to an edge from $E(H)$, which then can only lead outside of $K$. ∎
@@
Proof
If neither $w^{\prime}$ nor $w^{\prime\prime}$ are in $V(K^{\prime})$, there is nothing we have to show. Suppose $w^{\prime}\notin V(K^{\prime})$. Then also $w^{\prime\prime}\notin V(K^{\prime})$ since by Lemma 16, $w^{\prime}$ is mapped to $w^{\prime\prime}$ which by Invariant 3 would imply that both $w^{\prime}$ and $w^{\prime\prime}$ are in $V(K^{\prime})$. We therefore may assume $w^{\prime}\in V(K^{\prime})$. Now suppose that both $w^{\prime}$ and $w^{\prime\prime}$ are in $V(K^{\prime})$. Then there is a vertex $\tilde{w}\in V(K^{\prime})$ such that $\{w^{\prime},w^{\prime\prime},\tilde{w}\}$ form a triangle within $K^{\prime}$. By Invariant 4, the triangle is monochromatic and by Invariant 5, the edge in $H$ with color $\chi(w^{\prime}w^{\prime\prime})$ is contained in the triangle. However, this contradicts Lemma 16 according to which the edge in $H$ colored $\chi(w^{\prime}w^{\prime\prime})$ is contained in the triangle formed by $\{w,w^{\prime},w^{\prime\prime}\}$. Thus $w^{\prime\prime}\notin V(K^{\prime\prime})$.
Since we could not map $v$ to $w^{\prime\prime}$, by Lemma 16, $w^{\prime}$ is mapped to $w^{\prime\prime}$. If $w^{\prime}$ is an inner vertex, we have a contradiction since by Invariant 2, $w^{\prime}$ would be mapped to a vertex within $V(K^{\prime})$. Thus $w^{\prime}$ is an end vertex, which implies that via $w^{\prime}$, $v(K^{\prime})$ is mapped to $w^{\prime\prime}$. ∎
Then composing a new cactus from $K,K^{\prime}$, and the triangle formed by $\{w,w^{\prime},w^{\prime\prime}\}$ satisfies all conditions: Since by induction they are satisfied by $K$ and $K^{\prime}$, we only have to check the new triangle $\{w,w^{\prime},w^{\prime\prime}\}$. We obtain a plain cactus since the degrees of $w$ and $w^{\prime}$ are four and the degree of $w^{\prime\prime}$ is two. The inner vertex $v(K^{\prime})$ is mapped to $w^{\prime\prime}$ which implies Invariant 2 and 3. The triangle $\{w,w^{\prime},w^{\prime\prime}\}$ satisfies Invariant 4 by Lemma 16.
Since $ww^{\prime}\in E(H)$, Invariant 5 is satisfied, both $w$ and $w^{\prime}$ are in the scope of $Q$ and $K,K^{\prime}$ provide disjoint scope graphs, the conditions of Definition 5 and therefore Invariant 6 are satisfied. Invariant 7 follows from Invariant 6 and that $Q$ satisfies the conditions of Lemma 10. Finally, Invariant 8 follows since $w^{\prime\prime}$ is an end vertex and its incident edge from $H$ is not in the constructed cactus.
@@ 4 Subcubic Graphs and Claw-Free Graphs
Proof
Let $G$ be a normalized graph on $n$ vertices that has a perfect matching $M$. Recall that we can assume without loss of generality that $G$ is a connected graph with more than $2$ vertices. By Lemma 1, there is no feasible coloring of $G$ with more than $43n/$ 4 colors, and Algorithm 1 obtains at least $|M|=n/2$ colors. Thus the attained approximation ratio is at most $\frac{3n/4}{n/2}\leq 3/2$. ∎
@@
In order to handle subcubic graphs, we need to introduce additional modifications. The validity of these modifications follows from Lemma 6, which says that we can always assume that pendant edges are colored with a unique color.
@@
Figure 5: Modification 4: Bridge removal, when $\deg(u)=3$ and $\deg(v)=1$, and Modification 5, with $c_{1}$, $c_{2}$ and $c_{3}$ marked with red, blue and black edges, respectively.
Lemma 20
@@
Case 2: $\deg(u)=3$ and $\deg(v)=1$; the proof goes analogously if $\deg(u)=1$ and $\deg(v)=3$. Since we have a subcubic graph, $u$ has no neighbors besides $u_{1}$, $u_{2}$ and $v$. For every given optimal coloring $\chi^{\star}$ of $G$, we may assume that every pendant edge has a unique color, according to Lemma 6. Since $uv$ has a unique color $c_{1}$, the edges $u_{1}u$ and $u_{2}u$ have to be colored with the same color $c_{2}\neq c_{1}$.
@@
Case 3: $\deg(u)=\deg(v)=3$. Given an optimal coloring $\chi^{\star}$ we can assume that the edge $uv$ has a unique color due to the following argument. By removing the edge $uv$ from $G$ we obtain two graphs $G_{1},G_{2}$; let us consider the graphs $G_{1}^{+}:=G_{1}\cup\{uv\}$ and $G_{2}^{+}:=G_{2}\cup\{uv\}$. Notice that $uv$ is a pendant edge in both graphs. According to Lemma 6 we may assume that $G_{1}^{+}$ and $G_{2}^{+}$ have optimal colorings $\chi^{\star}_{1}$ and $\chi^{\star}_{2}$, respectively, such that $|\chi^{\star}_{i}(G_{i}^{+})|=|\chi^{\star}(G_{i}^{+})|$ holds for $i=1,2$ and $uv$ has a unique color in both: denote this color by $c_{1}$ and $c_{2}$, respectively. Now, identifying $c_{1}$ and $c_{2}$, taking the colors of $\chi^{\star}_{1}$ and $\chi^{\star}_{2}$ we can obtain a coloring for $G$, in which the bridge $uv$ has a unique color and which uses $|\chi^{\star}|$ many colors.
@@
As for the colorings of $G$ and $G^{\prime}$, we can use the arguments of Case 2. After the modification, all vertices of $G^{\prime}$ see the same colors as in $G$ and $|\chi^{\prime}|=|\chi^{\star}|$ holds and hence $\mathsf{opt}(G)\leq\mathsf{opt}(G^{\prime})$ is true. In the other direction, a coloring $\chi$ for $G$ can be obtained from a coloring $\chi^{\prime}$ for $G^{\prime}$, such that $|\chi|=|\chi^{\prime}|$. Therefore, Modification 4 is valid. ∎
@@
After applying Modifications 1–4, each component is therefore either a trivial component or a bridgeless cubic graph. It is well-known that each bridgeless cubic graph has a perfect matching [21]. Using the equivalence of modified graphs due to Lemma 2, the claim follows as a consequence of Lemma 19. ∎
@@
Proof
According to Lemma 6, we can assume that we have an optimal coloring $\chi^{\star}$ which assigns unique colors to pendant edges. Then $\chi^{\star}(u_{1}v_{1})=c_{1}$ and $\chi^{\star}(u_{2}v_{2})=c_{2}$, such that $c_{1}\neq c_{2}$ and no other edge is colored $c_{1}$ or $c_{2}$. But then $\chi^{\star}(u_{1}u_{2})=c_{3}$, where $c_{3}\notin\{c_{1},c_{2}\}$. Every other edge adjacent to $u_{1}$ or $u_{2}$ (besides $u_{1}v_{1}$ and $u_{2}v_{2}$) is colored $c_{3}$.
@@
In the other direction, suppose we have a graph $G^{\prime}$ with a pendant edge $u_{12}v_{12}$ at $u_{12}$, and an isolated edge $w_{1}w_{2}$. Using Lemma 6, for a coloring $\chi^{\prime}$ of $G^{\prime}$ we can construct a coloring $\hat{\chi}^{\prime}$ that assigns $u_{12}v_{12}$ a unique color $c_{1}$ and $|\chi^{\prime}|\leq|\hat{\chi}^{\prime}|$. Then, for each $v\neq v_{12}$, if the edge $u_{12}v$ exists it is colored $c_{3}\neq c_{1}$. Additionally, $\hat{\chi}^{\prime}(w_{1}w_{2})=c_{2}$ for some unique color $c_{2}$.
Let us now construct a graph $G$ by splitting vertex $u_{12}$ into two adjacent vertices $u_{1}$ and $u_{2}$, and replacing some vertices $u_{12}v$ with $u_{1}v$ or $u_{2}v$, or both, inheriting the color $c_{2}$ of $u_{12}v$. Furthermore, replace $u_{12}v_{12}$ by $u_{1}v_{1}$, and replace $w_{1}w_{2}$ by $u_{2}v_{2}$, and inherit the respective colors, $c_{1}$ and $c_{2}$. Finally, use the color $c_{3}$ for the edge $u_{1}u_{2}$. Observe that (regardless to possible equivalences between colors $c_{1}$, $c_{2}$ and $c_{3}$) this leads to a coloring $\chi$ feasible for $G$, with the same amount of colors than $\hat{\chi}^{\prime}$, so then $|\chi^{\prime}|\leq|\hat{\chi}^{\prime}|=|\chi|$ holds, finishing the proof. ∎
@@
Bounding the approximation factor. As a result, we arrive to a normalized graph $G^{\prime}$ with possibly more than one component. Consider first the case when $G^{\prime}$ is claw free; we argue about each component $C$ independently. Since $C$ is normalized, the optimal coloring of $C$ has at most $43|C|/$ 4 colors. If $|C|$ is even, it contains a perfect matching, which is of size $|C|/2$ due to Sumner [25] and Las Vergnas [18] and the claw-freeness of $G$. If $|C|$ is odd, we show that it contains a matching of size $(|C|-1)/2$. Remove a non-cutvertex $v$ from $C$ and denote the resulting connected graph $C^{\prime}$.
@@
Now let us consider the case when claws have been created during the normalization process. As shown in the first half of the proof, performing Modification 2 then Modification 5 results in creating a claw, but it also creates a new pendant edge and a new isolated edge in the process. Let us isolate the creation of one single claw, we will see that this argument can be repeated for all claws created. During Modification 2 and Modification 5 we remove a vertex from $G$ and add an isolated edge to it, which means the number of vertices is increased by $1$. On the other hand, we introduced $3$ new pendant vertices: $v_{12}$, $w_{1}$ and $w_{2}$. According to Lemma 1, the number of colors in a feasible coloring is therefore at most $43(n+1)/4-3/4=3n/$ 4. Observe that the size of the maximum matching did not change: it decreased by $1$ due to the contraction, but increased by $1$ due to the isolated edge; therefore Algorithm 1 still finds a coloring with at least $n/2$ colors (depending on the parity of $n$). Creating a claw this way thus always results in a graph, for which Algorithm 1 is a $3/2$-approximation. Since the pendant edge and the isolated edge do not get removed by any modification in a later step of the algorithm, we can simply repeat the same argument for every claw created. Thus, the claim follows. ∎
@@
Using the same arguments as in the proof of Lemma 3, we can assume that each triangle is monochromatic. Then, similarly to the single triangle case there is an edge in $G^{\prime}$ for each triangle in $G$ to allow for calculating a feasible coloring $\chi^{\prime}$ from $\chi$ of equal size, and vice versa. ∎
We remark that coming up with a perfect matching preserving modification that removes degree-$2$ vertices would yield a $1.5$-approximation for graphs that contain a perfect matching. Indeed, in that case the arguments in the proof of Lemma 19 would give us the result, as we could apply Modifications 1–3 to a graph containing a perfect matching until we obtain a normalized graph with a perfect matching. As we do not have a perfect matching preserving Modification 2, we use another approach, yielding a slightly worse approximation factor.
@@
Proof
Let $G$ be a graph that contains a perfect matching; let us fix such a matching $M$ for the rest of the proof. According to Lemma 22, Modification 1 and 3 do not violate the property of having a perfect matching. Therefore if we only perform these two modifications and arrive at a normalized graph, Algorithm 1 gives a $1.5$-approximation according to Lemma 19. We will now analyze how Modification 2 can affect the performance of Algorithm 1 using arguments similar to those in the case of claws in the proof of Theorem 1.2.
@@
Case 1: no pendant edge gets removed, therefore their number grows by $2$. Then the number of vertices is increased by $1$ and the upper bound is 3/$43/4(n+1)-1/2=3/4n+1/$ 4, by Lemma 1.
Case 2: one of the pendant edges gets removed, therefore their number grows by $1$. Then the number of vertices stays $n$, hence the upper bound is 3/$43/4n-1/$ 4. Notice that in this case we had a pendant edge $u_{1}w$ (or $u_{2}w$) in $G$, which was part of $M$ along with $u_{2}v$ (or $u_{1}v$), and the newly created $u_{1}v_{1}$ (or $u_{2}v_{2}$) got removed by Modification 1.

### v3

@@ 3.1 Preparing the character graph
Proof 3.22.
Since $K$ is a plain cactus and we cannot apply Modification $3$, there have to be edges not allowed in a simple cactus. These cannot be degree-$4$ vertices, due to Lemma 3.17. Also, by Lemma 3.19, edges between vertices from $K$ cannot violate the properties of simple cacti. Thus there has to be a degree-$2$ vertex $w$ of $K$ which has a set $A$ of at least two adjacent vertices in $V(G)\setminus V(K)$. By Invariant 8, there is a vertex $w^{\prime}\in A$ with $ww^{\prime}\in E(H)$. Let $c$ be the color of the triangle containing $w$ and $c^{\prime}:=\chi(ww^{\prime})$.
We claim that there is no vertex $\hat{w}\in A$ with $\chi(w\hat{w})\neq c^{\prime}$. Since $w$ only sees two colors, the only possibility would be $\chi(w\hat{w})=c$. However, since $w$ sees two colors, $\hat{w}$ cannot be an inner vertex and by Lemma 3.15 it also cannot be a free vertex or an end vertex.
Next, we show that $\iota(w^{\prime})=w^{\prime\prime}$. Suppose $\iota(w^{\prime})\neq w^{\prime\prime}$. Then, by Lemma 3.15, it sees two colors which are distinct from $c$, and also distinct from $c^{\prime}$ since $ww^{\prime}\in E(H)$ and $w,w^{\prime}\notin\iota^{-1}(w^{\prime\prime})$. However, $w^{\prime\prime}$ also sees $c^{\prime}$ and therefore more than two colors, a contradiction. Thus $\iota(w^{\prime})=w^{\prime\prime}$. In particular, $\chi(ww^{\prime})=\chi(ww^{\prime\prime})=\chi(w^{\prime}w^{\prime\prime})$.
With Lemma 3.21 we know that adding the triangle $\{w,w^{\prime},w^{\prime\prime}\}$ does not violate the conditions of plain cacti, i.e., adding it still satisfies Invariant 1. Since $\iota(w^{\prime})=w^{\prime\prime}$, also Invariants 2 and 3 are satisfied. Invariant 4 follows directly from Lemma 3.21. Since $ww^{\prime}\in E(H)$ by definition (within Algorithm 2), Invariant 5 follows. Invariant 6 follows by noting that $w\in\text{scope}(v)$, the edge $ww^{\prime}$ satisfies the conditions of the lemma, and thus the color class $c$ is added to the scope. Invariant 7 is a direct consequence of Invariant 6. Invariant 8 requires additional arguments.
Lemma 3.23.
@@
Proof 3.24.
Let $\tilde{u}$ be a degree-$2$ vertex. We distinguish the type of $\tilde{u}$. Due to Invariant 7, $\tilde{u}$ is not free. If $\tilde{u}$ is an inner vertex, only one of the two incident edges from $E(H)$ can be in $E(K)$, due to Invariant 4. Thus, due to the other edge, the invariant is satisfied. If $\tilde{u}$ is an end vertex, it has to have an incident edge $\tilde{u}u^{\prime}$ not in $E(K)$, due to Modification $2$. If $u^{\prime}\in V(K)$, Lemma 3.19 implies the invariant. We therefore may assume that $u^{\prime}\notin V(K)$.
By Invariant 3, there is a vertex $\bar{u}\in V(K)$ with $\iota(\bar{u})=\tilde{u}$. In particular, $\bar{u}$ is an inner vertex, $\bar{u}\tilde{u}\in E(K)$ and $\bar{u}\tilde{u}\notin E(H)$. Let $\hat{u}$ be the third vertex in the same triangle. By Invariant 4, $\chi(\hat{u}\tilde{u})=\chi(\hat{u}\bar{u})=\chi(\tilde{u}\bar{u})$. But then $\hat{u}\bar{u}\in E(H)$ as otherwise $\bar{u}$ would see three colors. Thus $\hat{u}\tilde{u}\notin E(H)$. However, since $u$ is an end vertex, it is incident to an edge from $E(H)$, which then can only lead outside of $K$.
@@
Proof 3.26.
If neither $w^{\prime}$ nor $w^{\prime\prime}$ are in $V(K^{\prime})$, there is nothing we have to show. Suppose $w^{\prime}\notin V(K^{\prime})$. Then also $w^{\prime\prime}\notin V(K^{\prime})$ since by Lemma 3.21, $w^{\prime}$ is mapped to $w^{\prime\prime}$ which by Invariant 3 would imply that both $w^{\prime}$ and $w^{\prime\prime}$ are in $V(K^{\prime})$. We therefore may assume $w^{\prime}\in V(K^{\prime})$. Now suppose that both $w^{\prime}$ and $w^{\prime\prime}$ are in $V(K^{\prime})$. Then there is a vertex $\tilde{w}\in V(K^{\prime})$ such that $\{w^{\prime},w^{\prime\prime},\tilde{w}\}$ form a triangle within $K^{\prime}$. By Invariant 4, the triangle is monochromatic and by Invariant 5, the edge in $H$ with color $\chi(w^{\prime}w^{\prime\prime})$ is contained in the triangle. However, this contradicts Lemma 3.21 according to which the edge in $H$ colored $\chi(w^{\prime}w^{\prime\prime})$ is contained in the triangle formed by $\{w,w^{\prime},w^{\prime\prime}\}$. Thus $w^{\prime\prime}\notin V(K^{\prime\prime})$.
Since we could not map $v$ to $w^{\prime\prime}$, by Lemma 3.21, $w^{\prime}$ is mapped to $w^{\prime\prime}$. If $w^{\prime}$ is an inner vertex, we have a contradiction since by Invariant 2, $w^{\prime}$ would be mapped to a vertex within $V(K^{\prime})$. Thus $w^{\prime}$ is an end vertex, which implies that via $w^{\prime}$, $v(K^{\prime})$ is mapped to $w^{\prime\prime}$.
Then composing a new cactus from $K,K^{\prime}$, and the triangle formed by $\{w,w^{\prime},w^{\prime\prime}\}$ satisfies all conditions: Since by induction they are satisfied by $K$ and $K^{\prime}$, we only have to check the new triangle $\{w,w^{\prime},w^{\prime\prime}\}$. We obtain a plain cactus since the degrees of $w$ and $w^{\prime}$ are four and the degree of $w^{\prime\prime}$ is two. The inner vertex $v(K^{\prime})$ is mapped to $w^{\prime\prime}$ which implies Invariant 2 and 3. The triangle $\{w,w^{\prime},w^{\prime\prime}\}$ satisfies Invariant 4 by Lemma 3.21.
Since $ww^{\prime}\in E(H)$, Invariant 5 is satisfied, both $w$ and $w^{\prime}$ are in the scope of $Q$ and $K,K^{\prime}$ provide disjoint scope graphs, the conditions of Definition 3.8 and therefore Invariant 6 are satisfied. Invariant 7 follows from Invariant 6 and that $Q$ satisfies the conditions of Lemma 3.9. Finally, Invariant 8 follows since $w^{\prime\prime}$ is an end vertex and its incident edge from $H$ is not in the constructed cactus.
@@ 4 Subcubic Graphs and Claw-Free Graphs
Proof 4.2.
Let $G$ be a normalized graph on $n$ vertices that has a perfect matching $M$. Recall that we can assume without loss of generality that $G$ is a connected graph with more than $2$ vertices. By Lemma 1.4, there is no feasible coloring of $G$ with more than $43n/$ 4 colors, and Algorithm 1 obtains at least $|M|=n/2$ colors. Thus the attained approximation ratio is at most $\frac{3n/4}{n/2}\leq 3/2$.
@@
In order to handle subcubic graphs, we need to introduce additional modifications. The validity of these modifications follows from Lemma 2.11, which says that we can always assume that pendant edges are colored with a unique color.
@@
Figure 5: Modification 4: Bridge removal, when $\deg(u)=3$ and $\deg(v)=1$.
Lemma 4.3.
@@
Case 2: $\deg(u)=3$ and $\deg(v)=1$; the proof goes analogously if $\deg(u)=1$ and $\deg(v)=3$. Since we have a subcubic graph, $u$ has no neighbors besides $u_{1}$, $u_{2}$ and $v$. For every given optimal coloring $\chi^{\star}$ of $G$, we may assume that every pendant edge has a unique color, according to Lemma 2.11. Since $uv$ has a unique color $c_{1}$, the edges $u_{1}u$ and $u_{2}u$ have to be colored with the same color $c_{2}\neq c_{1}$.
@@
Case 3: $\deg(u)=\deg(v)=3$. Given an optimal coloring $\chi^{\star}$ we can assume that the edge $uv$ has a unique color due to the following argument. By removing the edge $uv$ from $G$ we obtain two graphs $G_{1},G_{2}$; let us consider the graphs $G_{1}^{+}:=G_{1}\cup\{uv\}$ and $G_{2}^{+}:=G_{2}\cup\{uv\}$. Notice that $uv$ is a pendant edge in both graphs. According to Lemma 2.11 we may assume that $G_{1}^{+}$ and $G_{2}^{+}$ have optimal colorings $\chi^{\star}_{1}$ and $\chi^{\star}_{2}$, respectively, such that $|\chi^{\star}_{i}(G_{i}^{+})|=|\chi^{\star}(G_{i}^{+})|$ holds for $i=1,2$ and $uv$ has a unique color in both: denote this color by $c_{1}$ and $c_{2}$, respectively. Now, identifying $c_{1}$ and $c_{2}$, taking the colors of $\chi^{\star}_{1}$ and $\chi^{\star}_{2}$ we can obtain a coloring for $G$, in which the bridge $uv$ has a unique color and which uses $|\chi^{\star}|$ many colors.
@@
As for the colorings of $G$ and $G^{\prime}$, we can use the arguments of Case 2. After the modification, all vertices of $G^{\prime}$ see the same colors as in $G$ and $|\chi^{\prime}|=|\chi^{\star}|$ holds and hence $\mathsf{opt}(G)\leq\mathsf{opt}(G^{\prime})$ is true. In the other direction, a coloring $\chi$ for $G$ can be obtained from a coloring $\chi^{\prime}$ for $G^{\prime}$, such that $|\chi|=|\chi^{\prime}|$. Therefore, Modification 4 is valid.
@@
After applying Modifications 1–4, each component is therefore either a trivial component or a bridgeless cubic graph. It is well-known that each bridgeless cubic graph has a perfect matching [21]. Using the equivalence of modified graphs due to Lemma 2.3, the claim follows as a consequence of Lemma 4.1.
@@
Proof 4.7.
According to Lemma 2.11, we can assume that we have an optimal coloring $\chi^{\star}$ which assigns unique colors to pendant edges. Then $\chi^{\star}(u_{1}v_{1})=c_{1}$ and $\chi^{\star}(u_{2}v_{2})=c_{2}$, such that $c_{1}\neq c_{2}$ and no other edge is colored $c_{1}$ or $c_{2}$. But then $\chi^{\star}(u_{1}u_{2})=c_{3}$, where $c_{3}\notin\{c_{1},c_{2}\}$. Every other edge adjacent to $u_{1}$ or $u_{2}$ (besides $u_{1}v_{1}$ and $u_{2}v_{2}$) is colored $c_{3}$.
@@
In the other direction, suppose we have a graph $G^{\prime}$ with a pendant edge $u_{12}v_{12}$ at $u_{12}$, and an isolated edge $w_{1}w_{2}$. Using Lemma 2.11, for a coloring $\chi^{\prime}$ of $G^{\prime}$ we can construct a coloring $\hat{\chi}^{\prime}$ that assigns $u_{12}v_{12}$ a unique color $c_{1}$ and $|\chi^{\prime}|\leq|\hat{\chi}^{\prime}|$. Then, for each $v\neq v_{12}$, if the edge $u_{12}v$ exists it is colored $c_{3}\neq c_{1}$. Additionally, $\hat{\chi}^{\prime}(w_{1}w_{2})=c_{2}$ for some unique color $c_{2}$.
Let us now construct a graph $G$ by splitting vertex $u_{12}$ into two adjacent vertices $u_{1}$ and $u_{2}$, and replacing some vertices $u_{12}v$ with $u_{1}v$ or $u_{2}v$, or both, inheriting the color $c_{2}$ of $u_{12}v$. Furthermore, replace $u_{12}v_{12}$ by $u_{1}v_{1}$, and replace $w_{1}w_{2}$ by $u_{2}v_{2}$, and inherit the respective colors, $c_{1}$ and $c_{2}$. Finally, use the color $c_{3}$ for the edge $u_{1}u_{2}$. Observe that (regardless to possible equivalences between colors $c_{1}$, $c_{2}$ and $c_{3}$) this leads to a coloring $\chi$ feasible for $G$, with the same amount of colors than $\hat{\chi}^{\prime}$, so then $|\chi^{\prime}|\leq|\hat{\chi}^{\prime}|=|\chi|$ holds, finishing the proof.
@@
[…]
@@
Now let us consider the case when claws are being created during the normalization process. As shown in the first half of the proof, performing Modification 2 then Modification 5 results in creating a claw, but it also creates a new pendant edge and a new isolated edge in the process. Let us isolate the creation of one single claw, we will see that this argument can be repeated for all claws created. During Modification 2 and Modification 5 we remove a vertex from $G$ and add an isolated edge to it, which means the number of vertices is increased by $1$. On the other hand, we introduced $3$ new pendant vertices: $v_{12}$, $w_{1}$ and $w_{2}$. According to Lemma 1.4, the number of colors in a feasible coloring is therefore at most $43(n+1)/4-3/4=3n/$ 4. Observe that the size of the maximum matching did not change: it decreased by $1$ due to the contraction, but increased by $1$ due to the isolated edge; therefore Algorithm 1 still finds a coloring with at least $n/2$ colors (depending on the parity of $n$). Creating a claw this way thus always results in a graph, for which Algorithm 1 is a $3/2$-approximation. Since the pendant edge and the isolated edge do not get removed by any modification in a later step of the algorithm, we can simply repeat the same argument for every claw created. Thus, the claim follows.
@@
Using the same arguments as in the proof of Lemma 2.5, we can assume that each triangle is monochromatic. Then, similarly to the single triangle case, there is an edge in $G^{\prime}$ for each triangle in $G$ to allow for calculating a feasible coloring $\chi^{\prime}$ from $\chi$ of equal size, and vice versa.
We remark that coming up with a perfect matching preserving modification that removes degree-$2$ vertices would yield a $1.5$-approximation for graphs that contain a perfect matching. Indeed, in that case the arguments in the proof of Lemma 4.1 would give us the result, as we could apply Modifications 1–3 to a graph containing a perfect matching until we obtain a normalized graph with a perfect matching. As we do not have a perfect matching preserving Modification 2, we use another approach, yielding a slightly worse approximation factor.
@@
Proof 5.4.
Let $G$ be a graph that contains a perfect matching; let us fix such a matching $M$ for the rest of the proof. According to Lemma 5.2, Modification 1 and 3 do not violate the property of having a perfect matching. Therefore if we only perform these two modifications and arrive at a normalized graph, Algorithm 1 gives a $1.5$-approximation according to Lemma 4.1. We will now analyze how Modification 2 can affect the performance of Algorithm 1 using arguments similar to those in the case of claws in the proof of Theorem 1.2.
@@
Case 1: no pendant edge gets removed, therefore their number grows by $2$. Then the number of vertices is increased by $1$ and the upper bound is 3/4(n+1)1/2=3/$43/4(n+1)-1/2=3/4n+1/$ 4, by Lemma 1.4.
[…]